Evaluate: cot 30 degree- cot 75 degree/tan 15 degree -tan 60 degree

Grade:10

1 Answers

Arun
6 years ago
As we know
Cot 30 = cot (90-60)
But cot(90-60) = tan60
Similarily
Cot75 = cot (90-15) = tan15
Hence question becomes like below,
 
(tan 60 - tan15)/(tan15 - tan60)
= -1
Hope it helps